OpenStudy (anonymous):

write 3 equivalent fractions -4/4 -2/3 4/7

OpenStudy (anonymous):

-4/4=-2/2 -2/3=-6/9 4/7=8/14

OpenStudy (anonymous):

Do we need 3 for each one of them? or 3 in total?

OpenStudy (anonymous):

3 for each one

OpenStudy (anonymous):

@Jaspharmo